OpenStudy (staceyg):

-5/8x - 1/3y = 1/8. What would the graph look like for this equation?

OpenStudy (staceyg):

@tkhunny

OpenStudy (tkhunny):

That appears to be a linear equation. What's your favorite way to graph such a thing? I might be tempted to use the Intercept Form. Just multiply by 8. \(-5x - (8/3)y = 1\) That's not quite it. \(\dfrac{x}{-5} + \dfrac{y}{-3/8} = 1\) -- THERE it is! You may prefer the Slope-Intercept Form. Your choice.

OpenStudy (staceyg):

I'm using the slope-intercept form. I don't know how to graph it on though. Can you draw it?

OpenStudy (tkhunny):

It looks like any other line. Why is this one special? Show your Slope-Intercept Form.

OpenStudy (staceyg):

I mean I don't know how to graph it at all.